OpenStudy (anonymous):

Consider the attraction and repulsion of different types of charge. a) Explain why you expect to find that the amount of charge a battery can pump onto a parallel plate capacitor will double if the area of each plate doubles.

rvc (rvc):

@Mashy

OpenStudy (anonymous):

Do you already know how the capacitance could be related to the area?

OpenStudy (anonymous):

Well, the capacitance has a linear relation with the area of that plate making the capacitor. Now could you guess what the answer would become?

OpenStudy (anonymous):

is parallel plate equation C= (KeA)/(d) ?

OpenStudy (anonymous):

Yes, as we know \[C =k \epsilon_0A/d\] now, as it stated above we are going to make the area doubled, or we define the new area to be \[A'=2A\] substituting into the capacitance relation, we get\[C' =k \epsilon_0A'/d = k \epsilon_o2A/d=2(k \epsilon_0 A/d) = 2C\]